# Yuhong Liu, Yong Yang, Di Wang, Jie Jiang, Chunchao Guo

### v4

@@ Abstract
While 3D generative models have greatly improved artists’ workflows, the existing diffusion models for 3D generation suffer from slow generation and poor generalization. To address this issue, we propose a two-stage approach named  Hunyuan3D-1.0 including a lite version and a standard version, that both support text- and image-conditioned generation. In the first stage, we employ a multi-view diffusion model that efficiently generates multi-view RGB in approximately 4 seconds. These multi-view images capture rich details of the 3D asset from different viewpoints, relaxing the tasks from single-view to multi-view reconstruction. In the second stage, we introduce a feed-forward reconstruction model that rapidly and faithfully reconstructs the 3D asset given the generated multi-view images in approximately 7 seconds. The reconstruction network learns to handle noises and in-consistency introduced by the multi-view diffusion and leverages the available information from the condition image to efficiently recover the 3D structure. Our framework involves the text-to-image model, i.e., Hunyuan-DiT [22], making it a unified framework to support both text- and image-conditioned 3D generation. Our standard version has 3×3$\times$3 × more parameters than our lite and other existing model. Our  Hunyuan3D-1.0 achieves an impressive balance between speed and quality, significantly reducing generation time while maintaining the quality and diversity of the produced assets.
footnotetext: *Equal contribution.Corresponding author.
@@ 1 Introduction
Despite several works [59, 61, 40, 24, 1, 26, 27] in multi-view generation and sparse-view reconstruction, few have organized these approaches into a cohesive framework that addresses their combined challenges. First, widely used multi-view diffusion models are often criticized for multi-view inconsistency and slow denoising processes. Second, sparse-view reconstruction models typically rely solely on view-aware RGB images to predict 3D representations. Addressing these issues separately is challenging. Noticing the need to tackle these sub-tasks together, we propose Hunyuan3D-1.0, which integrates the strengths of multi-view diffusion models and sparse-view reconstruction models to achieve 3D generation in 10 seconds in the best-case scenario, achieving a subtle balance between generalization and quality. In the first stage, the multi-view diffusion model generates RGB to finish the 2D-to-3D lifting. We fine-tune a large-scale 2D diffusion model to generate multi-view images to enhance the model’s understanding of 3D information. Additionally, we set the 0-elevation camera orbit for the generated views to maximize the visible area between generated views. In the second stage, the sparse-view reconstruction model utilizes the imperfectly consistent multi-view images to recover the underlying 3D shape. Unlike most sparse-view reconstruction models that only use RGB images with known poses, we incorporate the conditional image, without the known view pose, to provide additional view information as an auxiliary input to cover the unseen part in the generated multi-view images. Furthermore, we employ a linear unpatchify layer operation to enrich details in the latent space without incurring additional memory or computational costs.
@@
  • We introduce a unified framework  Hunyuan3D-1.0, support text- and image- condition 3D generation both.
@@
Figure 1: The overview of our  Hunyuan3D-1.0. Given an input image, we first utilize a multi-view diffusion model to synthesize 6 novel views at fixed camera poses. Then we feed the generated multi-view images into a transformer-based sparse-view large reconstruction model to reconstruct a high-quality 3D mesh. The whole image-to-3D generation process takes only around 10 seconds.
@@ 3 Methods
We present the two stages in our approach, Hunyuan3D-1.0, in this section. First, we introduce the multi-view diffusion model for 2D-to-3D lifting in Sec. 3.1. Second, we discuss pose-known and pose-unknown image fusion and the super-resolution layer within the sparse-view reconstruction framework in Sec. 3.2.
@@ 3.2 Sparse-view Reconstruction Model
Figure 3: Qualitative comparisons of single-view generation. Our Hunyuan3D-1.0 achieves better visual quality compared to existing methods
@@
Table 1: Comparison on GSO [9]. Our Hunyuan3D-1.0 achieve new state-of-the-art performance on GSO [9] in terms of CD and F-score metrics.
@@
Table 2: Comparison on OminiObject3D [57]. Our Hunyuan3D-1.0 achieve new state-of-the-art performance on OmniObject3D [57] in terms of CD and F-score metrics.
Figure 4: User study. Our  Hunyuan3D-1.0 received the highest user preference across 5 metrics.
Figure 5: Performance vs Runtime. Our  Hunyuan3D-1.0 balance the quality and efficiency well.
@@ 5 Results
We quantitatively and qualitatively compare  Hunyuan3D-1.0 to previous state-of-the-art methods using two different datasets with 3D reconstruction metrics.
Quantitative Comparisons. We compare  Hunyuan3D-1.0 with the existing state-of-the-art baselines on 3D reconstruction that use feed-forward techniques, including OpenLRM [11], SyncDreamer [26], TripoSR [44], Wonder3D [27], CRM [53], LGM [42] and InstantMesh [59]. As shown in Table 1 and Table 3.2, our Hunyuan3D-1.0, especially our standard version, outperforms all the baselines, both in terms of CD and F-score metrics, achieving new state-of-the-art performance on this task.
Qualitative Comparisons. We present qualitative results of existing methods in Fig. 3. The figure illustrates that OperLRM [11] and TripoSR [44] struggle with geometric shapes, such as the soap and the box, and often generate blurred textures, as seen with the chair and the shoes. InstantMesh [20] captures more surface details but still exhibits some artifacts in certain areas, such as the seat of the chair, the logo on the cup, and the corners of the soap and box. In contrast, our model demonstrates superior reconstruction quality for both shape and texture. They not only capture the more accurate overall 3D structures of the objects but also excel in modeling intricate details. Our  Hunyuan3D-1.0 received the highest user preference across 5 metrics as shown in 4.
Performance vs. Runtime. Another key advantage of Hunyuan3D-1.0 is its inference speed. The lite model takes around 10 seconds to produce a 3D mesh from a single image, while the standard model takes roughly 25 seconds. Note that these times do not include UV map unwrapping and texture baking, which takes approximately 15 seconds. Fig. 5 presents a 2D plot comparing our method to existing approaches, with inference times on the x-axis and the average F-Score on the y-axis. The plot demonstrates that Hunyuan3D-1.0 achieves an optimal balance between quality and efficiency.
@@
This work introduces Hunyuan3D-1.0, a two-stage 3D generation pipeline capable of creating high-quality 3D shapes. The pipeline consists of a multi-view generation model that produces multi-view images rich in texture and geometry details and a feed-forward sparse-view reconstruction model that recovers the underlying 3D shape with explicit representations. We incorporate several innovative designs to enhance the speed and quality of the 3D generation process, including adaptive classifier-free guidance to balance the controllability and diversity for multi-view diffusion, hybrid inputs to address the unseen part reconstruction, and a lightweight super-resolution module to enhance the representation of details. Extensive evaluations on benchmark tasks demonstrate that Hunyuan3D-1.0 achieves state-of-the-art performance in 3D generation. Our method consistently outperforms existing approaches, highlighting its effectiveness in addressing the inherent challenges of 3D generation. These results validate the robustness and efficiency of our proposed pipeline, making substantial contributions to the 3D Generative community.

### v5

@@ Abstract
While 3D generative models have greatly improved artists’ workflows, the existing diffusion models for 3D generation suffer from slow generation and poor generalization. To address this issue, we propose a two-stage approach named  Hunyuan3D 1.0 including a lite version and a standard version, that both support text- and image-conditioned generation. In the first stage, we employ a multi-view diffusion model that efficiently generates multi-view RGB in approximately 4 seconds. These multi-view images capture rich details of the 3D asset from different viewpoints, relaxing the tasks from single-view to multi-view reconstruction. In the second stage, we introduce a feed-forward reconstruction model that rapidly and faithfully reconstructs the 3D asset given the generated multi-view images in approximately 7 seconds. The reconstruction network learns to handle noises and in-consistency introduced by the multi-view diffusion and leverages the available information from the condition image to efficiently recover the 3D structure. Our framework involves the text-to-image model, i.e., Hunyuan-DiT [22], making it a unified framework to support both text- and image-conditioned 3D generation. Our standard version has 3×3$\times$3 × more parameters than our lite and other existing model. Our  Hunyuan3D 1.0 achieves an impressive balance between speed and quality, significantly reducing generation time while maintaining the quality and diversity of the produced assets.
footnotetext: *Hunyuan3D team contributors are listed in the end of report.
@@ 1 Introduction
Despite several works [59, 61, 40, 24, 1, 26, 27] in multi-view generation and sparse-view reconstruction, few have organized these approaches into a cohesive framework that addresses their combined challenges. First, widely used multi-view diffusion models are often criticized for multi-view inconsistency and slow denoising processes. Second, sparse-view reconstruction models typically rely solely on view-aware RGB images to predict 3D representations. Addressing these issues separately is challenging. Noticing the need to tackle these sub-tasks together, we propose Hunyuan3D 1.0, which integrates the strengths of multi-view diffusion models and sparse-view reconstruction models to achieve 3D generation in 10 seconds in the best-case scenario, achieving a subtle balance between generalization and quality. In the first stage, the multi-view diffusion model generates RGB to finish the 2D-to-3D lifting. We fine-tune a large-scale 2D diffusion model to generate multi-view images to enhance the model’s understanding of 3D information. Additionally, we set the 0-elevation camera orbit for the generated views to maximize the visible area between generated views. In the second stage, the sparse-view reconstruction model utilizes the imperfectly consistent multi-view images to recover the underlying 3D shape. Unlike most sparse-view reconstruction models that only use RGB images with known poses, we incorporate the conditional image, without the known view pose, to provide additional view information as an auxiliary input to cover the unseen part in the generated multi-view images. Furthermore, we employ a linear unpatchify layer operation to enrich details in the latent space without incurring additional memory or computational costs.
@@
  • We introduce a unified framework  Hunyuan3D 1.0, support text- and image- condition 3D generation both.
@@
Figure 1: The overview of our  Hunyuan3D 1.0. Given an input image, we first utilize a multi-view diffusion model to synthesize 6 novel views at fixed camera poses. Then we feed the generated multi-view images into a transformer-based sparse-view large reconstruction model to reconstruct a high-quality 3D mesh. The whole image-to-3D generation process takes only around 10 seconds.
@@ 3 Methods
We present the two stages in our approach, Hunyuan3D 1.0, in this section. First, we introduce the multi-view diffusion model for 2D-to-3D lifting in Sec. 3.1. Second, we discuss pose-known and pose-unknown image fusion and the super-resolution layer within the sparse-view reconstruction framework in Sec. 3.2.
@@ 3.2 Sparse-view Reconstruction Model
Figure 3: Qualitative comparisons of single-view generation. Our Hunyuan3D 1.0 achieves better visual quality compared to existing methods
@@
Table 1: Comparison on GSO [9]. Our Hunyuan3D 1.0 achieve new state-of-the-art performance on GSO [9] in terms of CD and F-score metrics.
@@
Table 2: Comparison on OminiObject3D [57]. Our Hunyuan3D 1.0 achieve new state-of-the-art performance on OmniObject3D [57] in terms of CD and F-score metrics.
Figure 4: User study. Our  Hunyuan3D 1.0 received the highest user preference across 5 metrics.
Figure 5: Performance vs Runtime. Our  Hunyuan3D 1.0 balance the quality and efficiency well.
@@ 5 Results
We quantitatively and qualitatively compare  Hunyuan3D 1.0 to previous state-of-the-art methods using two different datasets with 3D reconstruction metrics.
Quantitative Comparisons. We compare  Hunyuan3D 1.0 with the existing state-of-the-art baselines on 3D reconstruction that use feed-forward techniques, including OpenLRM [11], SyncDreamer [26], TripoSR [44], Wonder3D [27], CRM [53], LGM [42] and InstantMesh [59]. As shown in Table 1 and Table 3.2, our Hunyuan3D 1.0, especially our standard version, outperforms all the baselines, both in terms of CD and F-score metrics, achieving new state-of-the-art performance on this task.
Qualitative Comparisons. We present qualitative results of existing methods in Fig. 3. The figure illustrates that OperLRM [11] and TripoSR [44] struggle with geometric shapes, such as the soap and the box, and often generate blurred textures, as seen with the chair and the shoes. InstantMesh [20] captures more surface details but still exhibits some artifacts in certain areas, such as the seat of the chair, the logo on the cup, and the corners of the soap and box. In contrast, our model demonstrates superior reconstruction quality for both shape and texture. They not only capture the more accurate overall 3D structures of the objects but also excel in modeling intricate details. Our  Hunyuan3D 1.0 received the highest user preference across 5 metrics as shown in 4.
Performance vs. Runtime. Another key advantage of Hunyuan3D 1.0 is its inference speed. The lite model takes around 10 seconds to produce a 3D mesh from a single image, while the standard model takes roughly 25 seconds. Note that these times do not include UV map unwrapping and texture baking, which takes approximately 15 seconds. Fig. 5 presents a 2D plot comparing our method to existing approaches, with inference times on the x-axis and the average F-Score on the y-axis. The plot demonstrates that Hunyuan3D 1.0 achieves an optimal balance between quality and efficiency.
@@
This work introduces Hunyuan3D 1.0, a two-stage 3D generation pipeline capable of creating high-quality 3D shapes. The pipeline consists of a multi-view generation model that produces multi-view images rich in texture and geometry details and a feed-forward sparse-view reconstruction model that recovers the underlying 3D shape with explicit representations. We incorporate several innovative designs to enhance the speed and quality of the 3D generation process, including adaptive classifier-free guidance to balance the controllability and diversity for multi-view diffusion, hybrid inputs to address the unseen part reconstruction, and a lightweight super-resolution module to enhance the representation of details. Extensive evaluations on benchmark tasks demonstrate that Hunyuan3D 1.0 achieves state-of-the-art performance in 3D generation. Our method consistently outperforms existing approaches, highlighting its effectiveness in addressing the inherent challenges of 3D generation. These results validate the robustness and efficiency of our proposed pipeline, making substantial contributions to the 3D Generative community.
8 Contributors
  • Project Sponsors: Jie Jiang, Yuhong Liu, Di Wang, Yong Yang, Tian Liu
  • Project Leaders: Chunchao Guo
  • Core Contributors:
    • Data: Lifu Wang, Jihong Zhang, Meng Chen, Liang Dong, Yiwen Jia, Yulin Cai, Jiaao Yu, Yixuan Tang, Hao Zhang, Zheng Ye, Peng He, Runzhou Wu, Chao Zhang, Yonghao Tan
    • Algorithm: Xianghui Yang, Huiwen Shi, Bowen Zhang, Fan Yang, Jiacheng Wang, Hongxu Zhao, Xinhai Liu, Xinzhou Wang, Qingxiang Lin, Jing Xu, Zebin He, Zhuo Chen, Sicong Liu, Junta Wu, Yihang Lian, Shaoxiong Yang
  • Contributors: Jie Xiao, Yangyu Tao, Jianchen Zhu, Jinbao Xue, Kai Liu, Chongqing Zhao, Xinming Wu, Zhichao Hu, Lei Qin, Jianbing Peng, Zhan Li, Minghui Chen, Xipeng Zhang, Lin Niu, Paige Wang, Yingkai Wang, Haozhao Kuang, Zhongyi Fan, Xu Zheng, Weihao Zhuang, YingPing He